OpenStudy (anonymous):

Let f(x) = 3x – 42 and g(x) = 5x. Part 1 What is (f ◦ g)(x) ? Part 2 Use complete sentences to describe the method you used to solve this problem.

OpenStudy (mertsj):

f composite g is the same as f(g(x)) do you know how to do that?

OpenStudy (mertsj):

f(g(x))=3(5x)-42=15x-42

OpenStudy (anonymous):

part 1 3(5x)-42 =15x-42 part 2 insert the value of g(x) in place of variable in f(x). solve!

OpenStudy (anonymous):

me gusta! thankies tiaraaaaa

OpenStudy (anonymous):

no problem! :)